# RouteNet-Gauss: Hardware-Enhanced Network Modeling with Machine Learning

### v1

@@ III-C RouteNet-Gauss’s architecture
3:$\boldsymbol{\hat{y}}^{(t)}_{f},\{\boldsymbol{h}^{(t+1)}_{q}|\forall q\in\mathcal{Q}\},\{\boldsymbol{h}^{(t+1)}_{d}|\forall d\in\mathcal{D}\}\leftarrow MP\big{(}\{\boldsymbol{h}^{(t)}_{f}|\forall f\in\mathcal{F}\},\{\boldsymbol{h}^{(t)}_{l}|\forall l\in\mathcal{L}\},\{\boldsymbol{h}^{(t)}_{q}|\forall q\in\mathcal{Q}\},\{\boldsymbol{h}^{(t)}_{d}|\forall d\in\mathcal{D}\}\big{)}$EndFor
@@
1:$pos\leftarrow 0$triangleright Curr. position in path
2:$\boldsymbol{h}^{t}_{f}\leftarrow\boldsymbol{h}^{t-1}_{f}$ForAll$(l,q)\in f$
@@
5:$M^{t}_{q}\leftarrow\boldsymbol{h}^{(t-1)}_{\hat{Q}_{d}(q)}\mathop{\parallel}\displaylimits\sum_{(f,pos)\in\hat{Q}_{f}(q)}{\widetilde{m}^{t}_{(f,pos)}}$
6:$\boldsymbol{h}^{t}_{q},\emptyset\leftarrow U_{Q}(M^{t}_{q};\boldsymbol{h}^{t-1}_{q})$
@@
10:$M^{t}_{lq}\leftarrow\widetilde{m}^{t}_{\hat{Q}_{q}(l)}$
11:$\boldsymbol{h}^{t}_{lq},\emptyset\leftarrow U_{L}(M^{t}_{lq};\boldsymbol{h}^{t-1}_{lq})$
@@
13:$\boldsymbol{\hat{y}}_{f}\leftarrow\sum_{\widetilde{m}^{t}_{f,pos}\in\widetilde{m}^{t}_{f}}R(\widetilde{m}^{t}_{f,pos})$EndFor
@@ V-B1 RouteNet-Gauss’s accuracy in synthetic traffic
TREX-S MAPE RouteNet-G 2.604% 2.790% 2.418% 2.591% 3.121%
@@
OMNeT++ $-4.337$ $-4.389$ $-3.741$ $-3.553$ $-3.184$
TREX-MB MAPE RouteNet-G 2.277% 2.480% 2.289% 2.529% 3.188%
@@
TREX-S MAPE RouteNet-G 9.447% 11.881% 9.347% 9.074% 9.609%
@@
OMNeT++ $-0.584$ $-0.509$ $-0.799$ $-0.702$ $-0.482$
TREX-MB MAPE RouteNet-G 10.711% 13.680% 10.934% 10.913% 11.824%
@@ V-B2 RouteNet-Gauss’s accuracy in real traffic data
Delay MAPE RouteNet-G 14.749% 12.372% 20.122% 24.316% 32.889%
OMNeT++ $93.075\%$ $82.611\%$ $131.792\%$ $147.495\%$ $158.056\%$
MAE ($\mu\text{s}$) RouteNet-G 47.217 27.522 134.366 187.471 289.536
OMNeT++ $126.088$ $90.462$ $283.835$ $368.629$ $503.644$
R2 RouteNet-G 0.055 0.045 0.011 -0.005 -0.037
OMNeT++ $-0.991$ $-1.017$ $-0.961$ $-0.934$ $-0.770$
Jitter MAPE RouteNet-G 14.445% 21.518% 14.994% 13.621% 16.493%
OMNeT++ $46.442\%$ $84.288\%$ $41.096\%$ $38.560\%$ $48.977\%$
MAE ($\mu\text{s}$) RouteNet-G 2.335 1.484 6.047 8.106 20.527
OMNeT++ $6.590$ $4.898$ $14.681$ $19.653$ $43.003$
R2 RouteNet-G 0.590 0.701 0.468 0.405 0.271
OMNeT++ $-0.959$ $-1.096$ $-0.912$ $-0.869$ $-0.765$
@@
Firstly, RouteNet-G continues to show better MAPE and MAE when predicting both the delay and jitter. For example, when predicting the delay RouteNet-G’s MAPE is up to $85\%$ lower than OMNeT++’s and its MAE is over 3.2 times smaller. Then, when predicting the jitter RouteNet-G’s MAPE is at least $67\%$ lower than OMNeT++’s. This is also validated by the illustration of the distributions, which show once again RouteNet-G’s delay distribution to be much closer to the ground truth. It should be noted that real traffic proves to be harder to model than synthetic ones as overall the three error metrics indicate poorer performance when compared to synthetic traffic.
Note that OMNeT++, as a DES-based simulator, is expected to obtain at least the same or better accuracy as ML models trained with simulated data (e.g. DQN, Parsimon). This suggests that a comparison between RouteNet-G and one of these alternatives would be even more favorable towards RouteNet-G. Overall, the results show that the benefits of RouteNet-G and using a testbed network also extend to captured traffic traces.
@@ VI-A2 Path lengths in larger networks
Another consideration in larger networks is the length of paths that flows follow. Since RouteNet-G treats paths as sequences of network elements (links and queues), it relies on seeing paths of similar lengths during training. This is mitigated by the fact that, in many computer networks, path lengths scale logarithmically with the number of nodes $N$ ($log(N)$ or even $\frac{log(N)}{log(log(N))}$[22]. The connectivity of the network and the routing protocol used also influence this behavior.

### v2

@@ III-C RouteNet-Gauss’s architecture
[…]
@@
[…]
@@ V-B1 RouteNet-Gauss’s accuracy in synthetic traffic
TREX-S
MAPE RouteNet-G 2.604% 2.790% 2.418% 2.591% 3.121%
@@
OMNeT++ 4.337-4.337 4.389-4.389 3.741-3.741 3.553-3.553 3.184-3.184
TREX-MB
MAPE RouteNet-G 2.277% 2.480% 2.289% 2.529% 3.188%
@@
TREX-S
MAPE RouteNet-G 9.447% 11.881% 9.347% 9.074% 9.609%
@@
OMNeT++ 0.584-0.584 0.509-0.509 0.799-0.799 0.702-0.702 0.482-0.482
TREX-MB
MAPE RouteNet-G 10.711% 13.680% 10.934% 10.913% 11.824%
@@ V-B2 RouteNet-Gauss’s accuracy in real traffic data
Delay
MAPE RouteNet-G 12.990% 8.904% 18.439% 21.638% $30.016\%$
OMNeT++ $55.808\%$ $64.239\%$ $39.786\%$ $34.860\%$ 28.182%
MAE ($\mu\text{s}$) RouteNet-G 44.564 24.515 $126.583$ $173.643$ $270.073$
OMNeT++ $57.350$ $55.541$ 62.892 68.590 91.492
R2 RouteNet-G $0.080$ $0.046$ $0.054$ $0.055$ $0.029$
OMNeT++ 0.939 0.931 0.972 0.952 0.811
Jitter
MAPE RouteNet-G $12.773\%$ $18.356\%$ $13.184\%$ $12.198\%$ $15.587\%$
RouteNet-F $14.938\%$ $22.531\%$ $15.418\%$ $14.142\%$ $20.768\%$
OMNeT++ 9.399% 16.056% 8.505% 8.816% 11.973%
MAE ($\mu\text{s}$) RouteNet-G $2.126$ $1.335$ $5.439$ $7.435$ $19.427$
RouteNet-F $2.371$ $1.591$ $6.018$ $8.036$ $21.424$
OMNeT++ 1.443 1.090 3.385 4.959 13.113
R2 RouteNet-G $0.628$ $0.735$ $0.513$ $0.437$ $0.320$
RouteNet-F $0.602$ $0.667$ $0.473$ $0.429$ $0.347$
OMNeT++ 0.843 0.840 0.825 0.788 0.662
@@
Overall, comparing higher error metrics when predicting real traffic to those when predicting synthetic traffic shows that the former is harder to predict. RouteNet-G and OMNeT++ perform similarly, with the specifics depending on the target performance metric. While predicting the delay, RouteNet-G outperforms DES when predicting the average and median delay, reducing the MAPE by $86\%$ and the MAE by $55.87\%$. When predicting the high percentile delays, both methods offer similar performance, with RouteNet-G scoring better MAPE values but worse MAE values. Overall, we also see RouteNet-G’s R2 scores are generally close to 0, while OMNeT++’s are high. Then, when predicting the jitter, OMNeT++ slightly yet consistently performs better than RouteNet-G: in terms of MAPE, for example, we see that OMNeT++ improves RouteNet-G’s score by up to $35\%$ (when predicting the 90th percentile jitter, from $13.184\%$ MAPE to $8.505\%$).
Altogether, the results show that RouteNet-G and OMNeT++ perform similarly, with differences mainly depending on the performance metric to predict. That being said, we must take into account that RouteNet-G can achieve these at a significantly lower computational cost. Note that OMNeT++, as a DES-based simulator, is expected to obtain at least the same or better accuracy as ML models trained with simulated data (e.g., DQN, Parsimon). This suggests that a comparison between RouteNet-G and one of these alternatives would be even more favorable towards RouteNet-G. Overall, the results show that the benefits of RouteNet-G and using a testbed network also extend to captured traffic traces.
@@ VI-A2 Path lengths in larger networks
Another consideration in larger networks is the length of paths that flows follow. Since RouteNet-G treats paths as sequences of network elements (links and queues), it relies on seeing paths of similar lengths during training. This is mitigated by the fact that, in many computer networks, path lengths scale logarithmically with the number of nodes NN ($log(N)$ or even $\frac{log(N)}{log(log(N))}$[22]. The connectivity of the network and the routing protocol used also influence this behavior.